Solve the IVP [] = [-3 1] [2+[6] [70] - [%]
Answer #1

(3)=441J (1) *=6911)h=644, 2-:] (94), ([@]=[J ze=certge 3 het z=Azth - Assim Pasticular Solutir Zp=dete zel=d z_arth de A (dz = zet ze General solution is z = get eget torte (1) ==)+63(3+3+() x(o)=4, 9(0)=-6 (1)= a (3)+cz(!) +0+ Card 4 = 26 € C2-15